# “Fear will keep the local systems in line.”

### v2

@@ 3.2. The case of $g=0$
Note that for a curve of genus 0, we have $\operatorname{LS}^{\operatorname{irred}}(X)=\emptyset$, so that (1.4) holds trivially.
@@ 3.3. The case $g=1$
In case (a), the dual group $\check{G}$} is isomorphic to $PGL_{n}$. In this case, [GLC4, Conjecture 4.5.7] is known (in fact, it is a trivial particular case of [BKS]). [^3] Fix an irreducible $PGL_{n}$ local system $\sigma$, and choose its generic lift to an $SL_{n}$-local system; denote the underlying vector bundle by $\mathcal{E}$}. Then the space of generic oper structures on $\sigma$ is isomorphic to the space of generically defined line subbundles in $\mathcal{E}$}, and this space is known to be homologically contractible by [Ga5].
@@ 4.2. Line bundles on $\operatorname{Bun}_{\check{G}}$
The fact that (4.4) is a perfect pairing implies that every $\mu_{\infty}$-torsor on every connected component of $\operatorname{Ge}_{\pi_{1}({\check{G}})}(X)$ is the restriction of ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$ for some ${\mathcal{P}}_{Z_{G}}$.
4.2.5.
@@
By Proposition 4.1.6, every étale local system of $k$-vector spaces on a given component of $\operatorname{Bun}_{\check{G}}$ splits as a direct sum of 1-dimensional ones, and each of the latter is isomorphic to the restriction of ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$ for some ${\mathcal{P}}_{Z_{G}}$.
4.2.6.
@@ 4.3. The fundamental group of $\operatorname{LS}_{\check{G}}^{\operatorname{irred}}$
Before we proceed to the proof, we record the following corollary, obtained by combining Theorem 4.3.2 and Proposition 4.1.6:
@@
Combining Propositions 4.3.5 and 4.1.6, we obtain:
Corollary 4.3.8.
Under the assumptions on $G$ and $g$ specified in Theorem 4.3.2, for every connected component of $\operatorname{Bun}_{\check{G}}^{\operatorname{stbl}}$, every étale local system of $k$-vector spaces on it splits as a direct sum of 1-dimensional ones. Each of the latter is isomorphic to the restriction of ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$ for some ${\mathcal{P}}_{Z_{G}}\in\operatorname{Bun}_{Z_{G}}$.
4.3.9.
@@
For the sake of completeness, we will supply a proof in Sect. 7.4. Combining Proposition 4.3.10 and Corollary 4.3.8, we obtain:
Corollary 4.3.11.
Under the assumptions on $G$ and $g$ specified in Theorem 4.3.2, for every connected component of $\operatorname{LS}_{\check{G}}^{\operatorname{stbl}}$, every étale local system of $k$-vector spaces on it splits as a direct sum of 1-dimensional ones. Each of the latter is isomorphic to the restriction of ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$ for some ${\mathcal{P}}_{Z_{G}}\in\operatorname{Bun}_{Z_{G}}$.
@@
Thus, given Corollary 4.3.11, in order to prove Theorem 4.3.2, it suffices to show that $\operatorname{LS}_{\check{G}}^{\operatorname{stbl}}$ is dense in $\operatorname{LS}_{\check{G}}^{\operatorname{irred}}$. This is equivalent to saying that $\operatorname{LS}_{\check{G}}^{\operatorname{stbl}}$ has a non-empty intersection with every irreducible component of $\operatorname{LS}_{\check{G}}^{\operatorname{irred}}$. Since $\operatorname{LS}_{\check{G}}^{\operatorname{irred}}$ is smooth, its irreducible components are the same as connected components.
@@
In it, the lower horizontal arrow and the left vertical arrow induce bijections on the sets of connected components, by Propositions 4.3.5 and 4.3.10, respectively.
@@ 5.1. Action of the center
$\pi_{1,\operatorname{alg}}({\check{G}})\simeq\pi_{0}(\operatorname{Bun}_{\check{G}})\simeq\pi_{0}(\operatorname{LS}_{\check{G}}).$
@@
We will prove (see Sect. 8.5.10):
@@
We will also prove (see Sect. 8.5.10):
@@ 5.4. Structure of ${\mathcal{A}}_{G,\operatorname{irred}}$
The restriction of ${\mathcal{A}}_{G,\operatorname{irred}}$ to every connected component is isomorphic to a direct sum of lines bundles, each of which is a restriction of some ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$ (see Sect. 4.2.6).
@@ 5.5.2. Step 1
$\widehat{\iota}^{!}({\mathcal{F}})\to{\mathcal{F}}\to\jmath\circ\jmath^{*}({\mathcal{F}}).$
@@
However, this follows from Proposition 5.3.5 and Corollary 5.3.3.
@@ 5.5.4. Step 2
Applying Step 1 (with ${\mathcal{L}}^{\otimes-1}_{{\mathcal{P}}_{Z_{G}}}$ instead of ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$, we find that
@@ 6.1. How to calculate those endomorphisms?
Let $r$ be the semi-simple rank of $G$. For each vertex of the Dynkin diagram we have a canonically defined map
$\operatorname{Bun}_{N,\rho(\omega_{X})}\to{\mathbb{A}}^{1}.$
The resulting map
$\operatorname{Bun}_{N,\rho(\omega_{X})}\to({\mathbb{A}}^{1})^{r}={\mathbb{A}}^{r}\overset{\operatorname{sum}}{\to}{\mathbb{A}}^{1}$
is the map $\chi^{\operatorname{glob}}$.
@@
$\chi^{\operatorname{glob}}/T:\operatorname{Bun}_{N,\rho(\omega_{X})}/T\to{\mathbb{A}}^{r}/T$
@@
Since the horizontal arrows in (6.3) are smooth with affine space fibers (and hence, the corresponding pullback functors are fully faithful), we obtain that in order to prove Theorems 5.2.3 and 5.2.4, it suffices to perform the corresponding calculation on ${\mathbb{A}}^{r}/T$.
@@
${\mathbb{A}}^{r}={\mathbb{A}}^{r}\times\operatorname{pt}\to{\mathbb{A}}^{r}\times\operatorname{pt}/Z_{G}\simeq{\mathbb{A}}^{r}/Z_{G}\to{\mathbb{A}}^{r}/T,$
and we have a Cartesian square
[…]
This reduces the assertions of Theorems 5.2.3 and 5.2.4 to the claim that the map
$k\to{\mathop{\operatorname{{\mathcal{E}}\it nd}}}({\mathsf{f}}^{\prime\prime}_{!}\circ\operatorname{sum}^{*}(\operatorname{exp}))$
is an isomorphism.
@@
Since
$\operatorname{sum}^{*}(\operatorname{exp})\simeq\operatorname{exp}^{\boxtimes r},$
by Künneth formula, we are reduced to showing that the map
$k\to{\mathop{\operatorname{{\mathcal{E}}\it nd}}}({\mathsf{f}}^{\prime\prime\prime}_{!}(\operatorname{exp}))$
is an isomorphism, where
${\mathsf{f}}^{\prime\prime\prime}:{\mathbb{A}}^{1}\to{\mathbb{A}}^{1}/{\mathbb{G}}_{m}.$
However, the latter is well-known: the object
${\mathsf{f}}^{\prime\prime\prime}_{!}(\operatorname{exp})\in\operatorname{D-mod}({\mathbb{A}}^{1}/{\mathbb{G}}_{m})$
is the *-direct image of $k\in\operatorname{D-mod}(\operatorname{pt})$ along
$\operatorname{pt}\simeq{\mathbb{A}}^{1}-0/{\mathbb{G}}_{m}\hookrightarrow{\mathbb{A}}^{1}/{\mathbb{G}}_{m}.$
@@ 7. Geometry of $\operatorname{Bun}_{G}$
The goal of this section is to prove Propositions 4.1.6, 4.3.5, 5.3.5 and 4.3.10.
For the duration of this section, we will change the notation from $\check{G}$} to $G$.
@@
such that
@@
(7.1) $\operatorname{ker}(\phi^{\prime})\simeq\pi_{1}(G).$

### v3

@@ 3.2. The case of $g=0$
Note that for a curve of genus 0, we have $\operatorname{LS}^{\operatorname{irred}}=\emptyset$, so that (1.4) holds trivially.
@@ 3.3. The case $g=1$
In case (a), the dual group $\check{G}$} is isomorphic to $PGL_{n}$. In this case, [GLC4, Conjecture 4.5.7] is known (in fact, it is a trivial particular case of [BKS]) [^8] Fix an irreducible $PGL_{n}$ local system $\sigma$, and choose its generic lift to an $SL_{n}$-local system; denote the underlying vector bundle by $\mathcal{E}$}. Then the space of generic oper structures on $\sigma$ is isomorphic to the space of generically defined line subbundles in $\mathcal{E}$}, and this space is known to be homologically contractible by [Ga5]..
@@ 4.2. Line bundles on $\operatorname{Bun}_{\check{G}}$
The fact that (4.4) is a perfect pairing implies:
Lemma 4.2.5.
Every $\mu_{\infty}$-torsor on every connected component of $\operatorname{Ge}_{\pi_{1}({\check{G}})}(X)$ is the restriction of ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$ for some ${\mathcal{P}}_{Z_{G}}$.
4.2.6.
@@
By Proposition 4.1.6, every étale local system of $k$-vector spaces on a given component of $\operatorname{Bun}_{\check{G}}$ splits as a direct sum of 1-dimensional ones, and by Lemma 4.2.5, each of the latter is isomorphic to the restriction of ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$ for some ${\mathcal{P}}_{Z_{G}}$.
4.2.7.
@@ 4.3. The fundamental group of $\operatorname{LS}_{\check{G}}^{\operatorname{irred}}$
Before we proceed to the proof, we record the following corollary, obtained by combining Theorem 4.3.2 and Proposition 4.1.6 (see Sect. 4.2.6):
@@
For the sake of completeness, we will supply a proof in Sect. 7.4.
@@
Next, we claim that $\operatorname{LS}_{\check{G}}^{\operatorname{stbl}}$ is dense in $\operatorname{LS}_{\check{G}}^{\operatorname{irred}}$. This is equivalent to saying that $\operatorname{LS}_{\check{G}}^{\operatorname{stbl}}$ has a non-empty intersection with every irreducible component of $\operatorname{LS}_{\check{G}}^{\operatorname{irred}}$. Since $\operatorname{LS}_{\check{G}}^{\operatorname{irred}}$ is smooth, its irreducible components are the same as connected components.
@@
In it, the lower horizontal arrow and the left vertical arrow induce bijections on the sets of connected components, by Propositions 4.3.5 and 4.3.8, respectively.
@@
Thus, we can work at one connected (=irreducible) component at a time (denote it by the subscript $\alpha$). We will show that the maps
(4.8) $\pi_{1}(\operatorname{LS}^{\operatorname{stbl}}_{{\check{G}},\alpha})\to\pi_{1}(\operatorname{LS}^{\operatorname{irred}}_{{\check{G}},\alpha})\to\pi_{1}(\operatorname{Bun}_{{\check{G}},\alpha})$
(for some choice of a base-point on $\operatorname{LS}^{\operatorname{stbl}}_{{\check{G}},\alpha}$) are isomorphisms.
The first map in (4.8) is surjective, thanks to the density. Hence, it is sufficient to show that the composite map is an isomorphism.
However, the above composite map is
$\pi_{1}(\operatorname{LS}^{\operatorname{stbl}}_{{\check{G}},\alpha})\to\pi_{1}(\operatorname{Bun}^{\operatorname{stbl}}_{{\check{G}},\alpha})\to\pi_{1}(\operatorname{Bun}_{{\check{G}},\alpha}).$
In the latter composition, the first map is an isomorphism thanks to Proposition 4.3.8. The corresponding fact for the second map follows from Proposition 4.3.5.
@@ 5.1. Action of the center
$\pi_{1,\operatorname{alg}}({\check{G}})\simeq\pi_{0}(\operatorname{Bun}_{\check{G}})\hookleftarrow\pi_{0}(\operatorname{LS}_{\check{G}}),$
and the latter inclusion is an isomorphism if $g\geq 2$.
@@
We will prove (see Sect. 8.6.10):
@@
We will also prove (see Sect. 8.6.10):
@@ 5.4. Structure of ${\mathcal{A}}_{G,\operatorname{irred}}$
The restriction of ${\mathcal{A}}_{G,\operatorname{irred}}$ to every connected component is isomorphic to a direct sum of lines bundles, each of which is a restriction of some ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$ (see Sect. 4.2.7).
@@ 5.5.2. Step 1
$\widehat{\iota}^{!}({\mathcal{F}})\to{\mathcal{F}}\to\jmath_{*}\circ\jmath^{*}({\mathcal{F}}).$
@@
However, this follows from Proposition 5.3.5 and Corollary 5.3.3:
If $Y$ is a Cohen-Macaulay scheme/stack and $i:Z\subset Y$ is a closed subscheme/substack of codimension $\geq k$, then, $\widehat{i}^{!}({\mathcal{O}}_{Y})$ lives in cohomological degrees $\geq k$.
@@ 5.5.4. Step 2
Applying Step 1 (with ${\mathcal{L}}^{\otimes-1}_{{\mathcal{P}}_{Z_{G}}}$ instead of ${\mathcal{L}}_{{\mathcal{P}}_{Z_{G}}}$), we find that
@@ 6.1. How to calculate those endomorphisms?
Let $r$ be the semi-simple rank of $G$. For each vertex $i$ of the Dynkin diagram $I$ we have a canonically defined map
$\chi^{\operatorname{glob},i}:\operatorname{Bun}_{N,\rho(\omega_{X})}\to{\mathbb{A}}^{1}.$
The composite map
$\operatorname{Bun}_{N,\rho(\omega_{X})}\overset{\chi^{\operatorname{glob},I}}{\to}({\mathbb{A}}^{1})^{r}={\mathbb{A}}^{r}\overset{\operatorname{sum}}{\to}{\mathbb{A}}^{1}$
is the map $\chi^{\operatorname{glob}}$ (see [GLC1, Sect. 1.3.6]).
@@
$\chi^{\operatorname{glob},I}/T:\operatorname{Bun}_{N,\rho(\omega_{X})}/T\to{\mathbb{A}}^{r}/T$
@@
Since the horizontal arrows in (6.3) are towers of torsors with respect to vector group-stacks [^13] I.e., we take group-stacks locally of the form ${\mathcal{E}}_{1}/{\mathcal{E}}_{0}$, where ${\mathcal{E}}_{0}$ and ${\mathcal{E}}_{1}$ are vector spaces, viewed as group-schemes. (and hence, the corresponding pullback functors are fully faithful), we obtain that in order to prove Theorems 5.2.3 and 5.2.4, it suffices to perform the corresponding calculation on ${\mathbb{A}}^{r}/T$.
I.e., we have to show that for every $\alpha\in(Z_{G})^{\vee}$, the map
$k\to{\mathop{\operatorname{{\mathcal{E}}\it nd}}}(({\mathsf{f}}^{\prime}_{!}\circ\operatorname{sum}^{*}(\operatorname{exp}))_{\alpha})$
is an isomorphism, where $({\mathsf{f}}^{\prime}_{!}\circ\operatorname{sum}^{*}(\operatorname{exp}))_{\alpha}$ is the corresponding direct summand of ${\mathsf{f}}^{\prime}_{!}\circ\operatorname{sum}^{*}(\operatorname{exp})$.
Note also that
(6.4) $\operatorname{sum}^{*}(\operatorname{exp})\simeq\operatorname{exp}^{\boxtimes r}.$
@@
${\mathbb{A}}^{r}={\mathbb{A}}^{r}\times\operatorname{pt}\to{\mathbb{A}}^{r}\times\operatorname{pt}/Z_{G}\simeq{\mathbb{A}}^{r}/Z_{G}\overset{{\mathsf{f}}^{\prime\prime}}{\to}{\mathbb{A}}^{r}/T.$
The direct summand $({\mathsf{f}}^{\prime}_{!}\circ\operatorname{sum}^{*}(\operatorname{exp}))_{\alpha}$ can be explicitly described as
(6.5) ${\mathsf{f}}^{\prime\prime}_{!}(\operatorname{sum}^{*}(\operatorname{exp})\boxtimes\psi_{\alpha}),$
where $\psi_{\alpha}$ is the (character) sheaf of $\operatorname{pt}/Z_{G}$ corresponding to the character $\alpha$.
@@
First, we claim that (6.5) is the *-extension of its own restriction to
$({\mathbb{A}}^{1}-0)^{r}/T\hookrightarrow{\mathbb{A}}^{r}/T.$
To prove this, we have to show this for each coordinate hyperplane separately, and it is sufficient to do so for the pullback of (6.5) to ${\mathbb{A}}^{r}$; denote it by ${\mathcal{G}}_{\alpha}$.
Thus, let us choose a vertex $i$ of the Dynkin diagram. Let
${\mathsf{k}}_{i}:{\mathbb{A}}^{r-1}\leftrightarrow{\mathbb{A}}^{r}:\pi_{i}$
be the inclusion and the projection into the corresponding hyperplane. We wish to show that ${\mathsf{k}}_{i}^{!}({\mathcal{G}}_{\alpha})=0$.
Choose a coweight of $T$ that is a positive multiple of the fundamental coweight corresponding to $i$ (the latter might itself not be a coweight of $T$). Note that the resulting action of ${\mathbb{G}}_{m}$ contracts ${\mathbb{A}}^{r}$ onto the above copy of ${\mathbb{A}}^{r-1}$.
Our object is equivariant with respect to $T$, and hence also with respect to the above action of ${\mathbb{G}}_{m}$. Hence, by the contraction principle,
${\mathsf{k}}_{i}^{!}({\mathcal{G}}_{\alpha})\simeq(\pi_{i})_{!}({\mathcal{G}}_{\alpha}).$
Now, $(\pi_{i})_{!}({\mathcal{G}}_{\alpha})$ is the pullback along ${\mathbb{A}}^{r-1}\to{\mathbb{A}}^{r-1}/T$ of
${\mathsf{f}}^{\prime\prime}_{!}((\pi_{i})_{!}(\operatorname{sum}^{*}(\operatorname{exp}))\boxtimes\psi_{\alpha}),$
where by a slight abuse of notation we denote by the same symbol $\mathsf{f}$$^{\prime\prime}$ the map
${\mathbb{A}}^{r-1}\times\operatorname{pt}/Z_{G}\overset{{\mathsf{f}}^{\prime\prime}}{\to}{\mathbb{A}}^{r-1}/T.$
However,
$(\pi_{i})_{!}(\operatorname{sum}^{*}(\operatorname{exp}))=0$
(this follows, e.g., from (6.4)).
6.2.6.
Thus, it remains to show that the (derived) endomorphisms of the restriction of (6.5) to the open substack $({\mathbb{A}}^{1}-0)^{r}/T$ are scalars.
Consider the corresponding map
$\overset{\circ}{{\mathsf{f}}}{}^{\prime\prime}:({\mathbb{A}}^{1}-0)^{r}\times\operatorname{pt}/Z_{G}\to({\mathbb{A}}^{1}-0)^{r}/T\simeq\operatorname{pt}/Z_{G},$
so that
${\mathsf{f}}^{\prime\prime}_{!}(\operatorname{sum}^{*}(\operatorname{exp})\boxtimes\psi_{\alpha})|_{({\mathbb{A}}^{1}-0)^{r}/T}\simeq(\overset{\circ}{{\mathsf{f}}}{}^{\prime\prime})_{!}(\operatorname{sum}^{*}(\operatorname{exp})|_{({\mathbb{A}}^{1}-0)^{r}}\boxtimes\psi_{\alpha}).$
Note that the map $\overset{\circ}{{\mathsf{f}}}{}^{\prime\prime}$ is equivariant with respect to $\operatorname{pt}/Z_{G}$. Hence, the object
(6.6) $(\overset{\circ}{{\mathsf{f}}}{}^{\prime\prime})_{!}(\operatorname{sum}^{*}(\operatorname{exp})|_{({\mathbb{A}}^{1}-0)^{r}}\boxtimes\psi_{\alpha})$
is a tensor multiple of $\psi_{\alpha}$.
It remains to show that this multiple is by a one-dimensional vector space. I.e., we have to show that the fiber of (6.6) at the point $1\in\operatorname{pt}/Z_{G}$ is 1-dimensional (in some cohomological degree).
6.2.7.
It is easy to see that the map $\overset{\circ}{{\mathsf{f}}}{}^{\prime\prime}$ equals the composition
$({\mathbb{A}}^{1}-0)^{r}\times\operatorname{pt}/Z_{G}\simeq{\mathbb{G}}_{m}^{r}\times\operatorname{pt}/Z_{G}\to\operatorname{pt}/Z_{G}\times\operatorname{pt}/Z_{G}\to\operatorname{pt}/Z_{G},$
where:
  • The second arrow is induced by the map ${\mathbb{G}}_{m}^{r}\simeq T_{\operatorname{ad}}\to\operatorname{pt}/Z_{G}$, corresponding to the cover $T\to T_{\operatorname{ad}}$;
  • The third arrow is the multiplication map.
From here we obtain that the fiber of (6.6) at $1\in\operatorname{pt}/Z_{G}$ identifies with
$\operatorname{C}^{\cdot}_{c}({\mathbb{G}}_{m}^{r},\operatorname{exp}^{\boxtimes r}|_{{\mathbb{G}}_{m}^{r}}\otimes\Psi^{\otimes-1}_{\alpha}),$
where $\Psi_{\alpha}$ is the Kummer sheaf on ${\mathbb{G}}_{m}^{r}$ corresponding to $\psi_{\alpha}$.
The question readily reduces to the case when $r=1$. I.e., we are interested in
(6.7) $\operatorname{C}^{\cdot}_{c}({\mathbb{G}}_{m},\operatorname{exp}|_{{\mathbb{G}}_{m}}\otimes\Psi),$
where $\Psi$ is a Kummer sheaf. We wish to show that this cohomology is 1-dimensional. This is well-known, but we supply an argument for completeness.
6.2.8.
We distinguish two scenarios.
If $\Psi$ is trivial, the result follows from the fiber sequence
$\operatorname{C}^{\cdot}_{c}({\mathbb{G}}_{m},\operatorname{exp}|_{{\mathbb{G}}_{m}})\to\operatorname{C}^{\cdot}_{c}({\mathbb{A}}^{1},\operatorname{exp})\to\operatorname{C}^{\cdot}_{c}(\operatorname{pt},\operatorname{exp}|_{0}),$
since the middle term vanishes.
If $\Psi$ is non-trivial, (6.7) is the fiber at the point $1$ in the dual ${\mathbb{A}}^{1}$ of the Fourier transform of (the clean) extension of $\Psi$ to ${\mathbb{A}}^{1}$. This Fourier transform is an irreducible perverse sheaf, which is equivariant for the action of ${\mathbb{G}}_{m}$ against (the inverse of) $\Psi$. Hence, this Fourier transform is a tensor multiple of (the inverse of) $\Psi$.
@@ 7. Geometry of $\operatorname{Bun}_{G}$
The goal of this section is to prove Propositions 4.1.6, 4.3.5, 5.3.5 and 4.3.8.
For the duration of this section, we will change the notation from $\check{G}$} to $G$, and we assume that it is semi-simple.
@@
such that [^14] Such a data is known as a z-extension in the sense of Kottwitz.
@@
(7.1) $\operatorname{ker}(\phi^{\prime})\simeq\pi_{1}(G):$
indeed
$\pi_{1}(G)\simeq\operatorname{ker}(G^{\prime}_{\operatorname{der}}\to G)\simeq\operatorname{ker}(T_{0}\to T_{1}).$